Gabi threw a 145 g baseball upwards with an initial speed of 25 m/s. How fast is the ball moving after 1 second?
1
Expert's answer
2020-12-14T12:12:51-0500

We can find the velocity of the ball after 1 second from the kinamatic equation:


v=v0gt=25 ms9.8 ms21 s=15.2 ms.v=v_0-gt=25\ \dfrac{m}{s}-9.8\ \dfrac{m}{s^2}\cdot 1\ s=15.2\ \dfrac{m}{s}.

Answer:

v=15.2 ms.v=15.2\ \dfrac{m}{s}.